OpenStudy (anonymous):

What is the fraction equivalent to 662/3%?

OpenStudy (anonymous):

66200/3 i believe

OpenStudy (anonymous):

how do you get the answer?

OpenStudy (anonymous):

@Pattywatty

OpenStudy (anonymous):

???????????????? @pattywatty

Directrix (directrix):

>66200/3 i believe ======== 662/3% Note: 3% = 3/100. Percent means hundredths. So, 662/ 3% = 662 / (3/100) = 662/1 * 100/3 = 66200/3 as Pattywatty posted above.